Suppose that X liters of 38% acid solution are mixed with y liters of a 42% solution to obtain 100L of a 39% solution. One equation in a system for solving this problem is x+y=100. Which one of the following is the other equation? A. 42x+38y=0.39*100 B 0.38x+0.42y=0.39*100 C. 0.42x+0.38y=0.39*100 D. 38x+42y=39
Answer by stanbon(75887)

Suppose that X liters of 38% acid solution are mixed with y liters of a 42% solution to obtain 100L of a 39% solution. One equation in a system for solving this problem is x+y=100.

Which one of the following is the other equation?
A. 42x+38y=0.39*100
B 0.38x+0.42y=0.39*100
C. 0.42x+0.38y=0.39*100
D. 38x+42y=39
=====
Ans: B
===========
Cheers,
Stan H.
===========